OpenStudy (anonymous):

What is the solution set of |x – 2| + 4 = 3? –3 or 3 –1 or 1 3 or 1 No solution

OpenStudy (nurali):

|x – 2| + 4 = 3 |x – 2| = 3-4 |x – 2| =-1 x-2=-(-1) or x – 2 = +(-1) x-2=1 or x-2=-1 x=3 or x=1 so the answer is 3 or 1